OpenStudy (anonymous):

how do you answer -2y=6(5-3x)

OpenStudy (anonymous):

Plz help

OpenStudy (shadowlegendx):

What are you solving for?

OpenStudy (anonymous):

slope intercept

OpenStudy (shadowlegendx):

Oh, change it into slope intercept form?

OpenStudy (anonymous):

yes

OpenStudy (shadowlegendx):

If so: -2y=6(5-3x) Use the Property of Distribution and distribute 6 to 5 and -3x -2y = 30 - 18x Divide both sides by -2 to isolate y y = -15 + 9x Rearrange problem to complete the form Slope Intercept Form: y = mx + b y = 9x - 15